OpenStudy (anonymous):

4cos^2(theta)+2sin(theta)-1=0

OpenStudy (anonymous):

this is much easier than the last one we can start by writing \[4(1-\sin^2(\theta))+2\sin(\theta)-1=0\] and then solve the quadratic in \(\sin(\theta)\)

OpenStudy (anonymous):

\[4-4\sin^2(\theta)+2\sin(\theta)-1=0\] \[4\sin^2(\theta)-2\sin(\theta)-3=0\] etc